OpenStudy (anonymous):

x-2y=3 4x-8y=12 solve?

hero (hero):

y = (3-x)/2 y = (12-4x)/-8 y = y -8(3-x) = 12-4x -24 + 8x = 12-4x 36 = 12x x = 3 y = 0

hero (hero):

Thanks for showing your appreciation for a proper solution.